# 3D and 4D World Modeling: A Survey

### v2

@@ 77 XCube [172] CVPR’24
[…]
@@ 88 SemCity [173] CVPR’24
[Uncaptioned image] [Uncaptioned image] N/A 1 LDM O3G, SSC $Uncaptioned image] \faExternalLink$
@@ 99 DriveWorld [174] CVPR’24
[…]
@@
[…]
@@
3 LDM O4G $Uncaptioned image] \faExternalLink$
@@
32 DiT O4G $Uncaptioned image] \faExternalLink$
@@
[…]
@@
[…]
@@
[…]
@@
- LDM O4G $Uncaptioned image] \faExternalLink$
@@
6 DiT O4G, OF $Uncaptioned image] \faExternalLink$
@@
[…]
@@
1 LDM O3G $Uncaptioned image] \faExternalLink$
@@ C   W   O
[…]
@@
[…]
$37$ T3Former [196] arXiv’25
N
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
T
[…]
$38$ COME [197] arXiv’25
N
[Uncaptioned image] [Uncaptioned image]
T
[…]
$39$ $\mathcal{X}$-Scene [198] arXiv’25
@@
1 LDM O3G $Uncaptioned image] \faExternalLink$
@@ 3.2 World Modeling from Occupancy Generation
Generation models based on occupancy grids tailored to offer a geometry-centric representation that encodes both semantic and structural details of the 3D world. By generating, forecasting, or simulating occupancy in 3D/4D space, these models provide a geometry-consistent scaffold for perception, enable action-contingent future prediction, and support realistic large-scale simulation. Based on their primary function, existing methods can be grouped into three categories: [Uncaptioned image] Scene Representors, [Uncaptioned image] Occupancy Forecasters, and [Uncaptioned image] Autoregressive Simulators. Table III summarizes existing models under these domains.
@@ 3.3 World Modeling from LiDAR Generation
LiDAR-based generation models provide geometry-aware and appearance-invariant representations by modeling complex scenes from point clouds. They enable robust 3D scene understanding and high-fidelity geometric simulation, offering advantages over image- and occupancy-based approaches in both geometric fidelity and environmental robustness. Based on their primary function, these methods can be classified into three categories: [Uncaptioned image] Data Engines, [Uncaptioned image] Action Interpreters, and [Uncaptioned image] Autoregressive Simulators. Table IV summarizes existing models under these domains.
@@ 3.3.1 Data Engines
Perception Data Augmentation. LiDAR-based generative modeling supports data augmentation for core 3D perception tasks such as detection and segmentation, with an emphasis on geometric fidelity and sensor realism. Early approaches primarily focused on modeling uncertainty and spatial structure to synthesize realistic LiDAR scans. DUSty [203] is a GAN-based framework that synthesizes realistic LiDAR scans by explicitly disentangling the underlying depth map from measurement uncertainty. DUSty v2 [204] extends DUSty by incorporating implicit neural representations, enabling the model to generate LiDAR range images at arbitrary resolutions. LiDARGen [205] pioneered the application of Langevin dynamics for LiDAR point cloud generation, achieving superior performance compared to GANs and VAEs. As the first work to adopt the denoising-diffusion paradigm in this domain, it has inspired numerous subsequent studies based on Denoising Diffusion Probabilistic Models (DDPMs) [68]. With explicit positional encoding, R2DM [206] achieves higher-precision LiDAR point cloud generation through a standardized DDPM process. Leveraging flow matching [207], R2Flow [208] significantly accelerates LiDAR point cloud generation. LiDM [209], RangeLDM [210], and 3DiSS [211] adopt latent diffusion technology by first compressing raw-scale data into low-dimensional latent variables through a pretrained VAE, then training the diffusion model in this latent space. The generated outputs are reconstructed to the original resolution, substantially improving generation speed while preserving quality. LiDARGRIT [212] extends this paradigm by discretizing the latent space with VQ-VAE [213] and generating latent codes using an autoregressive transformer. LiDARGRIT [212] further introduces a raydrop estimation loss to explicitly enhance the raydrop noise modeling. SDS [214] proposes simultaneous diffusion sampling for multi-view LiDAR scene generation, producing all views together to achieve much better geometric consistency than generating each view separately. Recently, SPIRAL [215] pioneered the generation of segmentation-labeled LiDAR data and introduced a novel closed-loop inference strategy that enhances consistency between geometry and semantics. La La LiDAR [22] proposes a layout-guided generative framework that integrates scene graph-based layout diffusion with a foreground-aware control injector, enabling explicit modeling of object relations and controllable scene generation. Veila [216] introduces a conditional diffusion framework for panoramic LiDAR generation guided by a monocular RGB image. It addresses the challenges of reliable conditioning, cross-modal alignment, and maintaining structural coherence beyond the RGB field of view. These advances enhance LiDAR-based perception by generating diverse, controllable, and geometrically faithful training data that capture real-world sensing characteristics.
TABLE IV: Summary of LiDAR-based generation (LiDARGen) models.
@@
$\bullet$ Datasets: K  KITTI [11], S  SemanticKITTI [16], N  nuScenes [10], 3  KITTI-360 [165], P  PandaSet [217] C  Carla [100], S  SeeingThroughFog [218], W  Waymo Open [95], N  NAVSIM [99], A  Argoverse 2 [97] and O  OmniDrive-nuScenes [219].
@@
$\bullet$ Categories: [Uncaptioned image] Data Engine (Sec. 3.3.1), [Uncaptioned image] Action Forecaster (Sec. 3.3.2), and [Uncaptioned image] Autoregressive Simulator (Sec. 3.3.3).
@@ 11 DUSty [203] IROS’21
[…]
@@ 22 LiDARGen [205] ECCV’22
[…]
@@ 55 Copilot4D [19] ICLR’24
6 LDM, AR L4G $Uncaptioned image] \faExternalLink$
@@ 66 R2DM [206] ICRA’24
[…]
@@ 77 ViDAR [221] CVPR’24
[Uncaptioned image] [Uncaptioned image] N/A 6 Enc-Dec L4G, DET, OP, E2E $Uncaptioned image] \faExternalLink$
@@ 99 LiDM [209] CVPR’24
[…]
@@
[…]
@@
6 LDM L4G, DET $Uncaptioned image] \faExternalLink$
@@
8 LDM L4G $Uncaptioned image] \faExternalLink$
@@
N/A LDM L4G, DET $Uncaptioned image] \faExternalLink$
@@
[…]
@@
N/A AR, LLM L4G, E2E $Uncaptioned image] \faExternalLink$
@@
[…]
$27$ 3DiSS [211] arXiv’25
S
  3
[…]
$28$ Distill-DPO [234] arXiv’25
S
[Uncaptioned image] [Uncaptioned image]
P
[…]
$29$ DriveX [235] arXiv’25
N
  N
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
T
6 Enc-Dec L4G, OP, E2E $Uncaptioned image] \faExternalLink$
$30$ OpenDWM [236] arXiv’25
@@ 3   N   W
N/A VQ-VAE LG, L4G $Uncaptioned image] \faExternalLink$
$31$ SPIRAL [215] arXiv’25
@@
[…]
$32$ La La LiDAR [22] arXiv’25
@@
[…]
$33$ Veila [216] arXiv’25
K
  S   N
[Uncaptioned image] [Uncaptioned image]
R
[…]
$34$ LiDARCrafter [49] arXiv’25
@@
66 Enc-Dec L4G $Uncaptioned image] \faExternalLink$
@@ 4 Datasets & Evaluations
TABLE V: Summary of datasets and benchmarks used for training VideoGen, OccGen, and LiDARGen models.
@@
$\bullet$ Tasked by: [Uncaptioned image] Video Generation Models (VideoGen, cf. Sec. 3.1), [Uncaptioned image] Occupancy Generation Models (OccGen, cf. Sec. 3.2), and [Uncaptioned image] LiDAR Generation Models (LiDARGen, cf. Sec. 3.3). Kindly refer to Table I for the definitions of conditions.
@@
$Uncaptioned image] \faExternalLink$
@@
$Uncaptioned image] \faExternalLink$
@@
[Uncaptioned image] $Uncaptioned image] \faExternalLink$
@@
[Uncaptioned image] $Uncaptioned image] \faExternalLink$
@@
$Uncaptioned image] \faExternalLink$
@@ 4.1 Datasets
In this survey, we discuss real, simulated, and augmented datasets that support research in 3D and 4D world modeling. These datasets span urban driving and related settings and provide rich annotations and conditions needed for VideoGen, OccGen, and LiDARGen. An overview of popular datasets and related benchmarks is illustrated in Figure 3. Additionally, Table V provides detailed statistics of each collection of the video, occupancy, LiDAR, and other relevant data formats from these mainstream datasets.
@@ 4.2 Evaluation Metrics & Protocols
Here, we organize evaluation metrics for world models into five perspectives. 1Generation Quality (Sec. 4.2.1) assesses the realism, coherence, and controllability of synthesized outputs. 2Forecasting Quality (Sec. 4.2.2) evaluates future predictions given partial observations. 3Planning-Centric Quality (Sec. 4.2.3) metrics measure safety and rule compliance in planning. 4Reconstruction-Centric Quality (Sec. 4.2.4) examines the ability of generation models to reproduce or simulate novel views. 5Downstream Evaluation (Sec. 4.2.5) tasks test how world models support tasks like detection, segmentation, and reasoning. A comprehensive summary of evaluation metrics is provided in Table LABEL:tab:summary_metrics. Together, these metrics cover both perceptual fidelity and utility in embodied decision-making and beyond.
@@ 4.2.1 Generation Quality
Generation quality focuses on whether a world model can produce realistic and coherent outputs given a prompt or condition. This involves four dimensions: fidelity, consistency, controllability, and human reference.
4.2.2 Forecasting Quality
4.2.3 Planning-Centric Quality
TABLE VI: Benchmarking VideoGen models on the Perceptual Fidelity of generation quality evaluations. The reported metrics are FID and FVD scores on the official nuScenes [10] validation set. All metrics are the lower the better ($\downarrow$).
@@ 4.2.3 Planning-Centric Quality
TABLE VII: Benchmarking VideoGen models on the Downstream Evaluation tasks. The reported metrics are mAP and NDS for 3D Object Detection, mIoU (Lanes, Drivable, Divider) for BEV Map Segmentation, L2 and Collision Rates at timestamps 1s, 2s, and 3s for Open-Loop Planning, and PDMS (P) and ADS (A) scores [119] for Closed-Loop Planning. All results are computed using the UniAD [254] implementation and checkpoints on the official nuScenes [10] validation set.
@@
TABLE VIII: Benchmarking VideoGen models on the Downstream Evaluation tasks. The reported metrics are mAP and NDS for 3D Object Detection (w/ BEVFusion [255] and StreamPETR [256]) and Road-wise mIoU scores (RmIoU) and Vehicle-wise mIoU scores (VmIoU) for BEV Map Segmentation (w/ CVT [257]). The results are on the official nuScenes [10] validation set. All metrics are the higher the better ($\uparrow$).

### v3

@@ 77 XCube [193] CVPR’24
[Uncaptioned image] [Uncaptioned image] N/A 1 LDM O3G [Uncaptioned image] W
@@ 88 SemCity [124] CVPR’24
[Uncaptioned image] [Uncaptioned image] N/A 1 LDM O3G, SSC [Uncaptioned image] [Uncaptioned image] W
@@ 4 MSSM OF, PT [Uncaptioned image] W
[Uncaptioned image] [Uncaptioned image] N/A 6 Enc-Dec OF [Uncaptioned image] W
@@
[Uncaptioned image] [Uncaptioned image] N/A 1 LDM O3G [Uncaptioned image] W
@@ 6 AR OF, E2E [Uncaptioned image] W
3 LDM O4G [Uncaptioned image] [Uncaptioned image] W
@@ 1 LDM O3G [Uncaptioned image] W
32 DiT O4G [Uncaptioned image] [Uncaptioned image] W
@@ 6 LLM O4G [Uncaptioned image] W
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] N/A 4 Enc-Dec OF, E2E [Uncaptioned image] W
@@ 11 DiT OF [Uncaptioned image] W
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] N/A 6 Enc-Dec OF, E2E [Uncaptioned image] W
@@
[Uncaptioned image] [Uncaptioned image] N/A 1 SD O3G [Uncaptioned image] W
@@ 6 LLM OF, E2E [Uncaptioned image] W
- LDM O4G [Uncaptioned image] [Uncaptioned image] W
@@ 4 Enc-Dec OF [Uncaptioned image] W
6 DiT O4G, OF [Uncaptioned image] [Uncaptioned image] W
@@
[Uncaptioned image] [Uncaptioned image] N/A 5 Enc-Dec OF [Uncaptioned image] W
@@
1 LDM O3G [Uncaptioned image] [Uncaptioned image] W
@@ C W O
[Uncaptioned image]/[Uncaptioned image] [Uncaptioned image] N/A 6 N/A OF [Uncaptioned image] W
@@
6 AR OF [Uncaptioned image] W
$37$ $\mathcal{X}$-Scene [280] NeurIPS’25
@@ 6 AR OF [Uncaptioned image] W
1 LDM O3G [Uncaptioned image] [Uncaptioned image] W
$38$ T3Former [261] arXiv’25
N
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
T
6 AR OF, E2E [Uncaptioned image] W
$39$ COME [204] arXiv’25
N
[Uncaptioned image] [Uncaptioned image]
T
6 DiT OF [Uncaptioned image] W
@@ 3.2 World Modeling from Occupancy Generation
Generation models based on occupancy grids tailored to offer a geometry-centric representation that encodes both semantic and structural details of the 3D world. By generating, forecasting, or simulating occupancy in 3D/4D space, these models provide a geometry-consistent scaffold for perception, enable action-contingent future prediction, and support realistic large-scale simulation. Based on their primary function, existing methods can be grouped into three categories: [Uncaptioned image] Scene Representors, [Uncaptioned image] Occupancy Forecasters, and [Uncaptioned image] Autoregressive Simulators. Table 3 summarizes existing models under these domains.
@@ 3.3 World Modeling from LiDAR Generation
LiDAR-based generation models provide geometry-aware and appearance-invariant representations by modeling complex scenes from point clouds. They enable robust 3D scene understanding and high-fidelity geometric simulation, offering advantages over image- and occupancy-based approaches in both geometric fidelity and environmental robustness. Based on their primary function, these methods can be classified into three categories: [Uncaptioned image] Data Engines, [Uncaptioned image] Action Interpreters, and [Uncaptioned image] Autoregressive Simulators. Table 4 summarizes existing models under these domains.
@@ 3.3.1 Data Engines
Perception Data Augmentation. LiDAR-based generative modeling supports data augmentation for core 3D perception tasks such as detection and segmentation, with an emphasis on geometric fidelity and sensor realism. Early approaches primarily focused on modeling uncertainty and spatial structure to synthesize realistic LiDAR scans. DUSty [174] is a GAN-based framework that synthesizes realistic LiDAR scans by explicitly disentangling the underlying depth map from measurement uncertainty. DUSty v2 [175] extends DUSty by incorporating implicit neural representations, enabling the model to generate LiDAR range images at arbitrary resolutions. LiDARGen [324] pioneered the application of Langevin dynamics for LiDAR point cloud generation, achieving superior performance compared to GANs and VAEs. As the first work to adopt the denoising-diffusion paradigm in this domain, it has inspired numerous subsequent studies based on Denoising Diffusion Probabilistic Models (DDPMs) [82]. With explicit positional encoding, R2DM [173] achieves higher-precision LiDAR point cloud generation through a standardized DDPM process. Leveraging flow matching [146], R2Flow [176] significantly accelerates LiDAR point cloud generation.
LiDM [192], RangeLDM [89], and 3DiSS [180] adopt latent diffusion technology by first compressing raw-scale data into low-dimensional latent variables through a pretrained VAE, then training the diffusion model in this latent space. The generated outputs are reconstructed to the original resolution, substantially improving generation speed while preserving quality. LiDARGRIT [78] extends this paradigm by discretizing the latent space with VQ-VAE [221] and generating latent codes using an autoregressive transformer. LiDARGRIT [78] further introduces a raydrop estimation loss to explicitly enhance the raydrop noise modeling. SDS [54] proposes simultaneous diffusion sampling for multi-view LiDAR scene generation, producing all views together to achieve much better geometric consistency than generating each view separately. Recently, SPIRAL [321] pioneered the generation of segmentation-labeled LiDAR data and introduced a novel closed-loop inference strategy that enhances consistency between geometry and semantics. La La LiDAR [148] proposes a layout-guided generative framework that integrates scene graph-based layout diffusion with a foreground-aware control injector, enabling explicit modeling of object relations and controllable scene generation. Veila [147] introduces a conditional diffusion framework for panoramic LiDAR generation guided by a monocular RGB image. It addresses the challenges of reliable conditioning, cross-modal alignment, and maintaining structural coherence beyond the RGB field of view. These advances enhance LiDAR-based perception by generating diverse, controllable, and geometrically faithful training data that capture real-world sensing characteristics.
Table 4: Summary of LiDAR-based generation (LiDARGen) models.
@@
$\bullet$ Datasets: K KITTI [69], S SemanticKITTI [14], N nuScenes [24], 3 KITTI-360 [141], P PandaSet [255] C Carla [48], S SeeingThroughFog [18], W Waymo [210], N NAVSIM [41], A Argoverse 2 [247] and O OmniDrive [232].
@@
$\bullet$ Categories: [Uncaptioned image] Data Engine (Sec. 3.3.1), [Uncaptioned image] Action Forecaster (Sec. 3.3.2), and [Uncaptioned image] AR Simulator (Sec. 3.3.3).
@@ 11 DUSty [174] IROS’21
[Uncaptioned image] [Uncaptioned image] N/A 1 GAN LG, SC [Uncaptioned image] W
@@ 22 LiDARGen [324] ECCV’22
[Uncaptioned image] [Uncaptioned image] N/A 1 Enc-Dec LG, SEG, SC [Uncaptioned image] W
@@ 55 Copilot4D [296] ICLR’24
6 LDM, AR L4G [Uncaptioned image] [Uncaptioned image] W
@@ 66 R2DM [173] ICRA’24
[Uncaptioned image] [Uncaptioned image] N/A 1 Enc-Dec LG, SC [Uncaptioned image] W
@@ 77 ViDAR [284] CVPR’24
[Uncaptioned image] [Uncaptioned image] N/A 6 Enc-Dec L4G, DET, OP, E2E [Uncaptioned image] [Uncaptioned image] W
@@ 1 LDM LG [Uncaptioned image] W
[Uncaptioned image] [Uncaptioned image] N/A 1 LDM LG, SC [Uncaptioned image] W
@@ 1 Enc-Dec LG, SC [Uncaptioned image] W
[Uncaptioned image] [Uncaptioned image] N/A 1 Enc-Dec, AR LG [Uncaptioned image] W
@@
6 LDM L4G, DET [Uncaptioned image] [Uncaptioned image] W
@@
8 LDM L4G [Uncaptioned image] [Uncaptioned image] W
@@ 1 LDM LG, DET [Uncaptioned image] W
N/A LDM L4G, DET [Uncaptioned image] [Uncaptioned image] W
@@ 1 Enc-Dec LG, SC, DET [Uncaptioned image] W
[Uncaptioned image] [Uncaptioned image] N/A 1 DiT LG [Uncaptioned image] W
@@ 1 Enc-Dec LG, SC [Uncaptioned image] W
N/A AR, LLM L4G, E2E [Uncaptioned image] [Uncaptioned image] W
@@
1 Enc-Dec LG, SC [Uncaptioned image] W
$27$ OpenDWM [200] CVPR’25
@@ 3 N W
N/A VQ-VAE LG, L4G [Uncaptioned image] [Uncaptioned image] W
$28$ SPIRAL [321] NeurIPS’25
@@
11 Enc-Dec LG, SEG [Uncaptioned image] W
$29$ 3DiSS [180] arXiv’25
S
3
[Uncaptioned image] [Uncaptioned image] N/A 1 LDM LG, SEG [Uncaptioned image] W
$30$ Distill-DPO [301] arXiv’25
S
[Uncaptioned image] [Uncaptioned image]
P
1 Enc-Dec LG, SC [Uncaptioned image] W
$31$ DriveX [203] arXiv’25
N
N
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
T
6 Enc-Dec L4G, OP, E2E [Uncaptioned image] [Uncaptioned image] W
$32$ Veila [147] arXiv’25
K
S N
[Uncaptioned image] [Uncaptioned image]
R
11 Enc-Dec LG, SEG [Uncaptioned image] W
$33$ La La LiDAR [148] AAAI’26
@@ 11 Enc-Dec LG, SEG [Uncaptioned image] W
11 Enc-Dec LG, SEG, DET, SC [Uncaptioned image] W
$34$ LiDARCrafter [139] AAAI’26
@@ 11 Enc-Dec LG, SEG, DET, SC [Uncaptioned image] W
66 Enc-Dec L4G [Uncaptioned image] [Uncaptioned image] W
@@ 4 Datasets & Evaluations
Table 5: Summary of datasets and benchmarks used for training VideoGen, OccGen, and LiDARGen models.
@@
$\bullet$ Tasked by: [Uncaptioned image] Video Gen. Models (VideoGen, cf. Sec. 3.1), [Uncaptioned image] Occupancy Gen. Models (OccGen, cf. Sec. 3.2), and [Uncaptioned image] LiDAR Gen. Models (LiDARGen, cf. Sec. 3.3). Kindly refer to Table 1 for the definitions of conditions.
@@
[Uncaptioned image] [Uncaptioned image] W
@@
[Uncaptioned image] [Uncaptioned image] W
@@
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] W
@@
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] W
@@
[Uncaptioned image] [Uncaptioned image] W
@@ 4.1 Datasets
In this survey, we discuss real, simulated, and augmented datasets that support research in 3D and 4D world modeling. These datasets span urban driving and related settings and provide rich annotations and conditions needed for VideoGen, OccGen, and LiDARGen. An overview of popular datasets and related benchmarks is illustrated in Figure 3. Additionally, Table 5 provides detailed statistics of each collection of the video, occupancy, LiDAR, and other relevant data formats from these mainstream datasets.
@@ 4.2 Evaluation Metrics & Protocols
Here, we organize evaluation metrics for world models into five perspectives:
  • 1Generation Quality (Sec. 4.2.1) assesses the realism, coherence, and controllability of synthesized outputs.
  • 2Forecasting Quality (Sec. 4.2.2) evaluates future predictions given partial observations.
  • 3Planning-Centric Quality (Sec. 4.2.3) metrics measure safety and rule compliance in planning.
  • 4Reconstruction-Centric Quality (Sec. 4.2.4) examines the ability of generation models to reproduce or simulate novel views.
  • 5Downstream Evaluation (Sec. 4.2.5) tasks test how world models support tasks like detection, segmentation, and reasoning.
A comprehensive summary of evaluation metrics is provided in Table LABEL:tab:summary_metrics. Together, these metrics cover both perceptual fidelity and utility in embodied decision-making and beyond.
@@ 4.2.1 Generation Quality
Table 6: Benchmarking VideoGen models on the Perceptual Fidelity of generation quality evaluations. The reported metrics are FID and FVD scores on the official nuScenes [24] validation set. All metrics are the lower the better ($\downarrow$).
@@
Generation quality focuses on whether a world model can produce realistic and coherent outputs given a prompt or condition. This involves four dimensions: fidelity, consistency, controllability, and human preference.
Table 7: Benchmarking VideoGen models on the Downstream Evaluation tasks. The reported metrics are mAP and NDS for 3D Object Detection, mIoU (Lanes, Drivable, Divider) for BEV Map Segmentation, L2 and Collision Rates at timestamps 1s, 2s, and 3s for Open-Loop Planning, and PDMS (P) and ADS (A) scores [278] for Closed-Loop Planning. All results are computed using the UniAD [92] implementation and checkpoints on official nuScenes [24] validation set.
@@
4.2.2 Forecasting Quality
Table 8: Benchmarking VideoGen models on the Downstream Evaluation tasks. The reported metrics are mAP and NDS for 3D Object Detection (w/ BEVFusion [151] and StreamPETR [231]) and Road-wise mIoU scores (RmIoU) and Vehicle-wise mIoU scores (VmIoU) for BEV Map Segmentation (w/ CVT [312]). The results are on the official nuScenes [24] validation set. All metrics are the higher the better ($\uparrow$).